Please can you help, this is my 10 year old daughter's homework!
I have tried lots of different possibilities.
x+a=750
x-a=34
what is the value of x?
what is the value of a?

Answer by Earlsdon(6294)

Find x and a from:
1) x+a = 750
2) x-a = 34 Add equations 1) and 2).
-----------
2x = 784 Divide both sides by 2.
x = 392
x+a = 750 Substitute x = 392.
392+a = 750 Subtract 392 from both sides.
a = 358